# Uncovering Singularities in Feynman Integrals via Machine Learning

### v1

@@ I Introduction
Around a decade ago, obtaining precise predictions for two-to-three particle collider processes beyond next-to-leading order was widely considered infeasible. This changed with advances in evaluating complicated two-loop Feynman integrals and interpreting them in terms of Chen iterated integrals. Key steps include deriving and solving differential equations for master integrals and assembling full amplitudes, often with finite-field techniques. In this context, the concept of the symbol alphabet and associated function spaces has become central for multi-loop studies Goncharov:2009lql ; Goncharov:2010jf . These tools capture the algebraic structure of iterated integrals, first explored by Chen in the 1970s Chen:1977oja , which naturally arise in canonical-form differential equations Henn:2013pwa and can be expressed as nested d-log integrals.
The symbol alphabet not only encodes the algebraic structure of iterated integrals but also underpins modern amplitude techniques, including bootstrap methods for master integrals and full scattering amplitudes Dixon:2011pw ; Dixon:2013eka ; Drummond:2014ffa ; Chicherin:2017dob ; Caron-Huot:2019vjl . Several tools facilitate symbol analysis: HyperInt implements a reduction algorithm yielding a superset of Landau singularities Panzer:2014caa , PLD.jl and SOFIA compute singularities Fevola:2023kaw ; Fevola:2023fzn ; Correia:2025yao , and Baikovletter attempts to reconstruct symbol alphabets via Baikov representations Jiang:2024eaj .
In this letter, we propose a symbolic-regression–based approach to uncover the singularity structure of multi-loop Feynman integrals. The framework combines efficient numerical evaluations with interpretable machine-learning techniques, enabling us to determine the canonical differential equations and to generate the complete symbol alphabet. Unlike previous machine-learning applications that focus mainly on accelerating IBP reductions vonHippel:2025okr ; Zeng:2025xbh ; Song:2025pwy , our method directly targets the analytic structure, requiring neither prior knowledge of singularities nor computationally costly reduction steps. We demonstrate the approach on nontrivial multi-loop examples, showing how all symbol letters—including square-root–type structures—can be systematically identified.
@@ II.1 From Feynman Integrals to Canonical Basis
where $v^{\mu}$ is an arbitrary vector built from the loop momenta or the external momenta. With the IBP identities in (2), a large linear system of Feynman integrals can be generated. By using Gaussian elimination, the basis of a given integral family can be found. The basis is called master integrals. The existence of master integrals is proved in Ref. Smirnov:2010hn .
@@
where the matrix $A$ encodes the analytic structure of the integrals. In the paper, we consider a broad class of Feynman integrals with polylogarithm expressions whose differential equation can be further decomposed as,
@@
with rational number matrices $A_{k}$ and algebraic kinematic functions $S_{k}$, known as symbol letters. The extension beyond algebraic letters is left for future investigation.
@@ III Symbolic Regression Frameworks and PySR
Symbolic regression aims to identify analytic expressions that capture relations between physical variables, with the functional form itself treated as an unknown. Mathematically, considering a grammar $\mathcal{G}$ of admissible operators (e.g., $\times,\sin,\exp$), variables, and constants, which generates a discrete space of symbolic structures $s\in\mathcal{S}$. Each structure admits continuous parameters $\theta\in\mathbb{R}^{k(s)}$, yielding functions $f_{s,\theta}:\mathbb{R}^{d}\to\mathbb{R}$. The task is thus a mixed discrete–continuous optimization: fitting continuous parameters within each structure while selecting the structure that achieves an optimal balance between accuracy and complexity. Formally, this can be written as
$(s^{\star},\theta^{\star})=\arg\min_{s\in\mathcal{S}}\left\{\min_{\theta\in\mathbb{R}^{k(s)}}\,\mathcal{L}_{D}(f_{s,\theta})+\lambda\,\mathcal{C}(s,\theta)\right\},$ (6)
where $\mathcal{L}_{D}$ is a data loss (e.g., the squared error with respect to physical measurements) and $\mathcal{C}(s,\theta)$ is a complexity penalty that encodes a preference for simple expressions. From a physics perspective, this separation reflects common practice: once a law is hypothesized in form, its constants are determined numerically. Moreover, symbolic regression naturally admits a multi-objective interpretation, in which predictive error $E(s,\theta)$ and model complexity $C(s)$ are treated as distinct optimization objectives. The Pareto front of non-dominated solutions provides a spectrum of candidate laws. It spans the simplicity–accuracy trade-off, from simple yet approximate to complex but precise, offering a principled balance between empirical adequacy and interpretability.
In practice, most SR frameworks rely on evolutionary search, where populations of candidate formulas are iteratively modified and selected for fitness schmidt2009distilling . Alternative approaches include reinforcement-learning strategies and neural methods (e.g. Deep Symbolic Regression and transformer-based decoders), which guide the search via learned policies or pre-trained large models biggio2021deep .
@@ III.1 PySR: A High-Performance Symbolic Regression Engine
PySR is engineered for speed and scalability. Operators are compiled into efficient kernels, evaluation uses JIT compilation, and populations are distributed across multiple CPU cores or clusters, enabling millions of expressions can be evaluated per second, outperforming pure-Python libraries such as gplearn. In addition, PySR provides a PySRRegressor class with a scikit-learn–like API. It outputs equations with associated loss and complexity, and models can be exported to SymPy or . By default, PySR returns the Pareto front of equations, allowing selection of candidate laws that balance accuracy and simplicity.
PySR combines distributed evolutionary search with efficient constant optimization, producing accurate and interpretable symbolic models. Its design makes it well-suited for methodological innovation and practical use in physics research, including applications to ODE/PDE discovery. Compared with other symbolic regression tools, PySR offers a favorable balance: Frameworks like gplearn are slower and less precise due to lack of constant optimization, while C++ libraries like Operon are extremely fast but less accessible and hard to extend. Neural-network-based approaches (e.g. AI-Feynmanudrescu2020ai ; udrescu2020ai2 ) can be powerful but require pre-training or large datasets. By contrast, PySR is both efficient and accessible, making it especially useful for physics applications where closed-form models are essential.
@@ III.2 Build Symbol Letters From Numerical Evaluations
Most symbolic regression methods aim to seek higher precision and interpretable symbolic expressions to describe meaningful relationships between physical observables and to predict results for unknown points in parameter space; And our work generates symbolic results through a ”forced” overfitting procedure, to obtain completely accurate symbolic expressions.
The workflow begins with the construction of a canonical basis for a given family of Feynman integrals. Numerical IBP reductions are performed with Kira Maierhofer:2017gsa ; Klappert:2020nbg ; Lange:2025fba to obtain CDE matrices at multiple kinematic points. (Since PySR performs regression with floating-point numbers, we truncate the rational coefficients from IBP reductions to 30 significant digits, balancing efficiency and accuracy.) Given up to $\mathcal{O}(100)$ pairs $(x_{i},\bm{A}(x_{i}))$, PySR searches for analytic candidates of the matrix elements and iteratively optimizes them. The resulting expressions are verified to exhibit a d-log form. Exponentiating and factorizing these entries yields a set of symbol letters, which constitute the candidate complete symbol alphabet for the integral family. The following figure 2 visualize our framework.
@@
For multi-variable cases, the ideas are similar but requires an additional modification of the template. As PySR does not have a complete template of multi-variable derivative, one has to define a template, containing not only variables $x_{1},x_{2},\dots$, but their derivatives $\partial f/\partial x_{1},\partial f/\partial x_{2},\dots$ in variable_names so that PySR will identify them as input features (see Listing B in Appendix B for complete code).
During the initial epochs, PySR may generate many candidate expressions with non-negligible errors (By default, PySR use Root Mean Square Error(RMSE) to estimate the regression error. initial errors may reach $10^{-2}\sim 10^{-5}$). After approximatly 40 epoches in less than one hour (depends on the complexity of target expressions), PySR identifies expressions formally equivalent to the target, with errors reduced to $\sim 10^{-30}$, consistent with the precision of the input. These symbolic expressions can then be exponentiated and factorized to extract the letters appearing in each entry.
@@ IV.2 Non-planar two-loop three-point Feynman integrals
$l_{3}=\sqrt{x+4},\quad l_{4}=\tfrac{1}{2}\!\left(\sqrt{x}+\sqrt{x-4}\right),\quad l_{5}=\sqrt{x-4}.$ (14)
@@ 5 Scales
[…]
   4        ——    ——    ——    ——
[…]
@@ Appendix B Options and Tuning of PySR Regressor
For multi-variable case, as demonstrated one should first declare a Hyper-parameter combine for PySR expression template:

### v2

@@ I Introduction
Around a decade ago, obtaining precise predictions for two-to-three particle collider processes beyond next-to-leading order was widely considered infeasible. This changed with advances in evaluating complicated two-loop Feynman integrals and interpreting them in terms of Chen’s iterated integrals. Key steps include deriving and solving differential equations for master integrals and assembling full amplitudes, often with finite-field techniques. In this context, the concept of the symbol alphabet and associated function spaces has become central for multi-loop studies Goncharov:2009lql ; Goncharov:2010jf . These tools capture the algebraic structure of iterated integrals, first explored by Chen in the 1970s Chen:1977oja , which naturally arise in canonical-form differential equations Henn:2013pwa and can be expressed as nested d-log integrals.
The symbol alphabet not only encodes the algebraic structure of iterated integrals but also underpins modern amplitude techniques, including bootstrap methods for master integrals and full scattering amplitudes Dixon:2011pw ; Dixon:2013eka ; Drummond:2014ffa ; Chicherin:2017dob ; Caron-Huot:2019vjl . Several tools facilitate symbol analysis: HyperInt implements a reduction algorithm yielding a superset of Landau singularities Panzer:2014caa , PLD.jl and SOFIA with Effortless implementation compute singularities Fevola:2023kaw ; Fevola:2023fzn ; Matijasic:2024too ; Caron-Huot:2024brh ; Correia:2025yao , and Baikovletter attempts to reconstruct symbol alphabets via Baikov representations Jiang:2024eaj .
In this letter, we propose a symbolic-regression–based approach to uncover the singularity structure of multi-loop Feynman integrals. The framework combines efficient numerical evaluations with interpretable machine-learning techniques, enabling us to determine the canonical differential equations and to generate the complete symbol alphabet. Unlike previous machine-learning applications that focus mainly on accelerating IBP reductions vonHippel:2025okr ; Zeng:2025xbh ; Song:2025pwy , our method directly targets the analytic structure, requiring neither prior knowledge of singularities nor computationally costly analytical reduction steps. We demonstrate the approach on nontrivial multi-loop examples, showing how all symbol letters—including square-root–type structures—can be systematically identified.
@@ II.1 From Feynman Integrals to Canonical Basis
where $v^{\mu}$ is an arbitrary vector built from the loop momenta or the external momenta. With the IBP identities in (2), a large linear system of Feynman integrals can be generated. By Gaussian elimination, the basis of a given integral family can be found. The basis is called master integrals. The existence of master integrals is proved in Ref. Smirnov:2010hn .
@@
where the matrix $A$ encodes the analytic structure of the integrals. In the letter, we consider a broad class of Feynman integrals with polylogarithm expressions whose differential equation can be further decomposed as,
@@
with rational number matrices $A_{k}$ and algebraic kinematic functions $S_{k}$, known as symbol letters. This covers a wide range of cases that appear in calculations from the Standard Model and gravitational-wave physics, and the extension beyond algebraic letters is left for future investigation.
@@ III Symbolic Regression Frameworks and PySR
Symbolic regression aims to identify analytic expressions that capture relations between physical variables, with the functional form itself treated as an unknown. Mathematically, considering a grammar $\mathcal{G}$ of admissible operators (e.g., $\times,\sin,\exp$), variables, and constants, which generates a discrete space of symbolic structures $s\in\mathcal{S}$. Each structure admits continuous parameters $\theta\in\mathbb{R}^{k(s)}$, yielding functions $f_{s,\theta}:\mathbb{R}^{d}\to\mathbb{R}$. The task is thus a mixed discrete–continuous optimization: fitting continuous parameters $\theta$ within each structure while selecting the structure $s$ that achieves an optimal balance between accuracy and complexity. Formally, this can be written as
$(s^{\star},\theta^{\star})=\arg\min_{s\in\mathcal{S}}\left\{\min_{\theta\in\mathbb{R}^{k(s)}}\,\mathcal{L}_{D}(f_{s,\theta})+\lambda\,\mathcal{C}(s)\right\},$ (6)
where $\mathcal{L}_{D}$ is a data loss (e.g., the squared predictive error with respect to physical measurements) and $\mathcal{C}(s,\theta)$ is a complexity penalty that encodes a preference for simple expressions. From a physics perspective, this separation reflects common practice: once a law is hypothesized in form, its constants are determined numerically. Moreover, symbolic regression naturally admits a multi-objective interpretation, in which predictive error $E(s,\theta)$ and model complexity $C(s)$ are treated as distinct optimization objectives. The Pareto front of non-dominated solutions provides a spectrum of candidate laws. It spans the simplicity–accuracy trade-off, from simple yet approximate to complex but precise, offering a principled balance between empirical adequacy and interpretability.
In practice, most SR frameworks rely on evolutionary search, where populations of candidate formulas are iteratively modified and selected for fitness schmidt2009distilling ; Song:2025odk . Alternative approaches include reinforcement-learning strategies and neural methods (e.g. Deep Symbolic Regression and transformer-based decoders), which guide the search via learned policies or pre-trained large models biggio2021deep .
@@ III.1 PySR: A High-Performance Symbolic Regression Engine
PySR combines distributed evolutionary search with efficient constant optimization, producing accurate and interpretable symbolic models. Its design makes it well-suited for methodological innovation and practical use in physics research, including applications to ODE/PDE discovery. Compared with other symbolic regression tools, PySR offers a favorable balance: Frameworks like gplearn are slower and less precise due to lack of constant optimization, while C++ libraries like Operon10.1145/3377929.3398099 are extremely fast but less accessible and hard to extend. Neural-network-based approaches (e.g. AI-Feynmanudrescu2020ai ; udrescu2020ai2 ) can be powerful but require pre-training or large datasets. By contrast, PySR is both efficient and accessible, making it especially useful for physics applications where closed-form models are essential.
@@ III.2 Build Symbol Letters From Numerical Evaluations
Most symbolic regression methods aim to seek higher precision and interpretable symbolic expressions to describe meaningful relationships between physical observables and to predict results for unknown points in parameter space. These symbolic relations are not necessary to be absolutely correct. In our work, however, we generates symbolic results through a ”overfitting-directed” tuning procedure, to obtain completely accurate symbolic expressions.
The workflow begins with the construction of a canonical basis for a given family of Feynman integrals. Numerical IBP reductions are performed with Kira Maierhofer:2017gsa ; Klappert:2020nbg ; Lange:2025fba to obtain CDE matrices at multiple kinematic points. (Since PySR performs regression with floating-point numbers, we truncate the rational coefficients from IBP reductions to 30 significant digits, balancing efficiency and accuracy.) Given up to $\mathcal{O}(100)$ pairs $(x_{i},{A}(x_{i}))$, PySR searches for analytic candidates of the matrix elements and iteratively optimizes them. The resulting expressions are verified to exhibit a d-log form. Exponentiating and factorizing these entries yields a set of symbol letters, which constitute the candidate complete symbol alphabet for the integral family. The following figure 2 visualizes our framework.
@@
During the initial epochs, PySR may generate many candidate expressions with non-negligible errors (By default, PySR use Root Mean Square Error(RMSE) to estimate the regression error. initial errors may reach $10^{-1}\sim 10^{-3}$). One should carefully tune Hyper-parameters to avoid underfitting. After approximatly 40 epoches in less than one hour (depends on the complexity of target expressions), PySR identifies expressions formally equivalent to the target, with errors reduced to $\sim 10^{-30}$, consistent with the precision of the input. This means PySR has find a proper expression equivalent to the correct answer, and corresponding symbolic expressions can be exponentiated and factorized to extract the letters appearing in each entry. One can now stop the searching to avoid overfitting and extract symbol letters from the results, and this can be automated by tuning Hyper-parameter early_stop_condition.
@@ IV.2 Non-planar two-loop three-point Feynman integrals
$l_{3}=\sqrt{x+4},\quad l_{4}=\tfrac{1}{2}\!\left(\sqrt{x}+\sqrt{x-4}\right),\quad l_{5}=\sqrt{x-4}.$
@@ 5 Scales
   1    ✓✓    ✓✓    ✓✓
   2    ✓✓    ✓✓    ✓✓
   3    ✓✓    ✓✓    ✓✓        ——
   4    ✓✓    ——    ——    ——    ——
Table 2: Trials on multi-loop, multi-scale integral families. ✓✓: complete symbol alphabet reconstructed; : all even letters obtained, remaining ones (odd letters) can be constructed manually; : some letters not found; ——: not tested because of incomplete reference. An $x$-p $y$-m integral has $x$ external momenta, of which $y$ are massive.
@@ Appendix B Options and Tuning of PySR Regressor
For multi-variable case, this is not a standard usage of PySR. To achieve this, one should first declare a Hyper-parameter combine for PySR expression template: